Solve Proportions: 9/-15 = -27/P
My question is: If just the demonitor or just the numerator is a negative number does this make the entire fraction negative?
Thank you for your help,
Brooke

Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

9/-15=-9/15=-(9/15)
-9/15=-27/P NOW CROSS MULTIPLY.
-9P=15*-27
-9P=-405
P=-405/-9
P=45 ANSWER.

Answer by stanbon(75887)


Yes.
Every fraction has three signs: sign of the fraction, sign of the numerator,
sign of the denominator. You may change any two of the signs and still have
the same fraction.
-------------------------------
9/-15 = -27/P
Invert both sides to get:
-15/9 = P/-27
Reduce to get:
-5/3 = P/-27
Multiply both sides by -27 to solve for P:
P = -27(-5/3)
P = 45
================
Cheers,
Stan H.